Bob is a W-2 employee for ABC Company. In December of 2019, Bob receives a promotion and his employer tells him that his new compensation will be $160,000 in 2020. Bob’s 2019 compensation was $130,000. He asks you to calculate how much his annual take home pay will increase for 2020. Bob does not participate in any retirement or health plans through his employer. He withholds 15% of his gross pay for federal income taxes and 5% of his gross pay for state income taxes.
Answer #1

Bob's 2019 gross pay was $ 130,000 and deductions were 15% towards federal and 5% for state taxes, totalling 20% from gross pay

Deductions = $ 130,000 × 20% = $ 26,000

Net pay for 2019 = Gross pay - Deduction = $ 130,000 - 26000

Net pay for 2019 = $ 104,000.

For 2020, his gross pay increase to $ 160,000 and dedication are 20% ,

So deduction for 2020 = $ 160,000 × 20% = $ 32,000

Net pay for 2020= $ 160,000 - 32000

Net pay for 2020 = $ 128,000.

So Increase of Net pay from 2019 to 2020 = Net pay of 20 - Net pay of 2019 = $ 128,000 - $ 104,000

Annual take home pay will increase for 2020 = $ 24,000.

If taken as a % = 24000 / 104000 = 23.08%.

Feel free to comment doubts if any